OpenStudy (anonymous):

whats the reference angle for 14π/3?

OpenStudy (mathstudent55):

14(pi)/3 - 2(pi) = 8(pi)/3 8(pi)/3 - 2(pi) = 2(pi)/3

OpenStudy (anonymous):

thanks :) so i just simplify from there?

OpenStudy (anonymous):

would 2.09 be correct? @mathstudent55

Directrix (directrix):

From @mathstudent55 above, 2(pi)/3 is the reference angle. I don't usually approximate but you can check your answer by entering "2 times pi divided by 3=" in the Google search query section.

OpenStudy (anonymous):

oh ok :) i'm confused though. when it asks for the reference angle should i give the number that i get when i do 2(pi)/3 on the calculator, or should i just enter 2(pi)/3 as the final answer?

OpenStudy (mathstudent55):

2(pi)/3 is the reference angle. That's the exact value of it in radians.

OpenStudy (anonymous):

oh okay great, thanks so much!! :)

OpenStudy (mathstudent55):

wlcm